1. A skater of mass 8000 g is carrying a box of mass 5 kg. The skater has a speed of 6 m/s with respect to the floor and is gliding without any friction on a smooth surface. a. Find the momentum of the box with respect to the floor. b. Find the momentum of the box with respect to the floor after she puts the box down on the frictionless skating surface.
  2. A 650 g hammer is moving horizontally at 9.00 m/s when it strikes a nail and comes to rest after driving the nail 5.00 cm into a board. Assume constant acceleration of the hammer-nail pair. a. Calculate the duration of the impact. b. What was the average force exerted on the nail?
1
Expert's answer
2021-07-30T03:39:04-0400

1.The momentum of the box with respect to the floor

"p=mv=5\:\rm kg\times 6\: m/s=30\: kg\cdot m/s"

2. The duration of the impact

"t=\frac{2d}{v}=\frac{2*0.05\: \rm m}{9.00\:\rm m/s}=0.011\:\rm s"

The average force exerted on the nail

"F=\frac{mv}{t}=\frac{0.650\;\rm kg*9.00\: m/s}{0.011\:\rm s}=527\:\rm N"